A steel cylinder has a volume of 25.0 L and is filled with N2 gas to a pressure of 200 atm. @ 30.0°C. How many moles of N2 gas does the cylinder contain? 



1
Expert's answer
2022-03-25T12:30:01-0400

30.0oC = 273.15 + 30.0 = 303.15 K


"n=\frac{PV}{RT}=\frac{200\ atm\times25.0\ L}{0.08206\frac{L\cdot{atm}}{mol\cdot{K}}\times303.15\ K}=201\ mol"


Answer: 201 mol